Consider the market for oranges. The supply curve is QS=-20 +10P The demand curve is Qd...

  1. Consider the market for oranges.

The supply curve is QS=-20 +10P

The demand curve is Qd =100-10P

  1. Find the Equilibrium price.
  2. Find the equilibrium quantity
  3. Draw a rough sketch of your curves and depict the equilibrium
  4. What will be the outcome if the government fixes the price at $5.00 what will be the outcome?
  5. Calculate the consumers’ and producers’ surplus at the equilibrium price.
Answer #1

a. At equilibrium Qs = ad -20 + lop = 100 -10P - 200 = 120 P26 Q t = 100-10 (6) = 40 b. = 2 d. At p=5, Qd = 100-10 (5)=so - Q